6. Calculate the empirical formula: C=58.5% H=4.1% N=11.4% and 0-26.0%.
Answer #1

Element   Mass   Number of moles    Division by small number   Nearest whole number

C             58.5   58.5/12 = 4.875         4.875 / 0.814 = 5.99                           6

H              4.1     4.1 / 1 = 4.1                4.1 / 0.814 = 5.04                             5

N            11.4    11.4 / 14 = 0.814        0.814 / 0.814 = 1                               1

O           26.0     26.0 / 16 = 1.625         1.625 / 0.814 = 1.996                       2

Therefore,

The emperical formula = C6H5NO2